Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 397     Entry time: 20 Oct 2023, 23:10
Author: Romain Letrun 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 13.6 keV, ~1.6 mJ and up to 352 pulses at 1.1 MHz.
  • Beam down at 22:47, not recovered before the end of the shift. Otherwise stable delivery.

Optical laser delivery

  • 800 nm, 50 fs, 1 to 176 pulses at 564 kHz pumping TOPAS at 475 nm.
  • Intensity modulations through the burst during the second half of the day. Solved by LAS OCD. Otherwise stable delivery.

Achievements/Observations

  • Remounting of diagnostics stack under the injector.
  • PP laser and X-ray spatial and timing overlap.
  • Collected data from noon until beam down towards the end of the shift.

Issues

  • Discrepancies between delay calculated by OpticalDelay device and motor encoder position appearing from time to time.
  • Karabo GUI crashed twice when after pressing enter while editing a table.
ELOG V3.1.4-7c3fd00